A box in the shape of a cube has 12-inch edges. What is the surface area of the box? show work?

Answer by mananth(16946)

the box is 12 inch side cube
how many surfaces are there to the box ?
There are 6 surface to a box.
the area of each surface = 12 *12 = 144 square inches
there are 6 sides
so total surface area of the box = 6 * 144 = 864 square inches